OpenStudy (flagged):

What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1

jimthompson5910 (jim_thompson5910):

Since we want integer coefficients, we should get those decimal values to whole numbers. To do this, we multiply both sides by 10 y = -6.9x + 5.1 10*y = 10*(-6.9x + 5.1) 10y = 10*(-6.9x) + 10*(5.1) 10y = -69x + 51 Are you following so far?

OpenStudy (waterlover9150802):

Since you have y=-6.9x+5.1 and standard form is Ax+By=C. You have to add 6.9x so you would have 6.9x+y=5.1